# Near-horizon polarized images of rotating hairy Horndeski black holes

### v1

@@ III Impacts of the Hairy Parameter on the Near-Horizon Polarized Image of a rotating Horndeski Black Hole
Let us now to discuss features of the near-horizon polarized image of a rotating Horndeski black hole (1) and to probe the corresponding impacts of the hairy parameter $h$. In Fig.(3), we present variations of the ratio $z$ with the quantity $\Delta$ in the rotating hairy Horndeski black hole spacetime for three different plasma flow conditions as in Hou:2023bep . The plasma is assumed to infall from infinity in the type I model and to plunge from the prograde innermost stable circular orbit (ISCO) in the type II model, respectively. The type III model corresponds to the fitting result from the simulation of MAD introduced in Chael:2021rjo . Our results show that the curves $z-\Delta$ are almost overlapped in the range $\Delta<10^{-2}$ for all three types of plasma flows in the black hole spacetimewith different values of the hairy parameter $h$. This further confirms that effects arising from types of plasma flow on the polarization structures near black hole horizon are negligible. Furthermore, from their partial enlargement, we find that the curves $z-\Delta$ are separate for different hairy parameter $h$, while still are overlapped for above three types of plasma flows. This means that the effects from the hairy parameter $h$ is much stronger than those from the types of plasma flows. In additional, the ratio $z$ increases with the hairy parameter $h$.
@@
Figure 4: Variation of the EVPAs in the near horizon images with the azimuthal angle $\phi$ in the rotating hairy Horndeski black hole spacetime. The left panel is for $\theta_{o}=17^{\circ}$ and the right one is for $\theta=80^{\circ}$. In each panel, the solid and dashed lines correspond to the cases $h=h_{c}$ and $h=0$, respectively. The blue and red lines denote the cases $a=0.94$ and $a=0.2$, respectively.
@@
In Fig.4, we present the changes of the observed EVPA of polarization vectors with the the azimuth angle $\phi$ for the photons from region near the event horizon for a rotating hairy Horndeski black hole with different hairy parameter $h$. As expected, the observed EVPA of polarization vectors significantly changes with the black hole’s spin $a$ and the observed inclination $\theta_{o}$. Furthermore, we find that the effects of hairy parameter $h$ on the observed EVPA $\chi$ depend on the black hole’s spin $a$ and the observed inclination $\theta_{o}$. For a fixed inclination $\theta_{o}$, in the slowly rotating case where $a=0.2$, the observed EVPA $\chi$ corresponding to $h=h_{c}$ is greater than that associated with $h=0$. The situation is reversed in the rapidly rotating case with $a=0.94$. For a fixed black hole spin, as the azimuthal angle $\phi$ is in the range $(3,6)$, one can find that the impact of $h$ on the observed EVPA is more distinct in the low inclination case with $\theta_{o}=17^{\circ}$. As the azimuthal angle $\phi$ is beyond the range $(3,6)$, the difference between the observed EVPAs induced by the hairy parameter $h$ is almost independent of the observer inclination $\theta_{o}$. Therefore, the observer inclination’s impact on the EVPA difference caused by the hairy parameter $h$ is also shaped by the azimuthal angle of the observed polarization vector. Through making a comparison, We also find that the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin $a$ than to the inclination angle $\theta_{o}$.
@@
Figure 5: The EVPAs in the near horizon images of rotating hairy Horndeski black holes are represented by different symbols: $\bullet$ indicates $\chi=\frac{\pi}{6}$, and $\triangle$ indicates $\chi=-\frac{\pi}{6}$. The left panel is for $a=0.2$ and the right one is for $a=0.94$. In each panel, the curves represent the near horizon images. The solid and dashed lines correspond to the cases $h=h_{c}$ and $h=0$, respectively. The blue and red lines denote the cases $\theta_{o}=17^{\circ}$ and $\theta_{o}=80^{\circ}$, respectively.
@@
In Fig.5, we present the positions of the polarization vectors related to EVPA $\chi=\frac{\pi}{6}$ or $\chi=-\frac{\pi}{6}$ in the near horizon images for a rotating hairy Horndeski black hole with spin parameters $a=0.2$ and $a=0.94$, observed at inclination angles $\theta_{o}=17^{\circ}$ and $\theta_{o}=80^{\circ}$. It is show that the position of polarization vectors with the same EVPA $\chi=\frac{\pi}{6}$ or $\chi=-\frac{\pi}{6}$ changes with the hairy parameter $h$. Apart from the size of the event horizon image, the variation in the position of polarization vectors with the same EVPA, induced by the hairy parameter $h$, is most strongly affected by the black hole spin. Although the effects from the hairy parameter $h$ in the near horizon images is weaker than those from black hole spin and observed inclination, one can infer the information on black hole’s hair by integrating the corresponding observations with higher precision.
@@ IV Discussions and conclusions
We have investigated the observed polarization in the near-horizon images of a rotating hair black hole within Horndeski gravity. Besides its mass and spin parameter, this black hole possesses an additional hairy parameter $h$, which originates from Horndeski theory. Our results confirm once again that the EVPA of polarization vectors in the near-horizon images is governed exclusively by spacetime geometry and is irrespective of the plasma flow geometry as found in Hou:2024qqo . Furthermore, the effects of the hairy parameter $h$ on the observed EVPA depend on the black hole’s spin $a$ and the observed inclination $\theta_{o}$. For a fixed inclination $\theta_{o}$, in the slowly rotating case where $a=0.2$, the observed EVPA corresponding to $h=h_{c}$ is greater than that associated with $h=0$. The situation is reversed in the rapidly rotating case with $a=0.94$. For a fixed black hole spin, the dependence of the difference between the observed EVPAs induced by the hairy parameter $h$ on the observer inclination $\theta_{o}$ is also influenced by the azimuthal angle of the observed polarization vector. As the azimuthal angle $\phi$ is in the range $(3,6)$, the impact of $h$ on the observed EVPA is more distinct in the low inclination case with $\theta_{o}=17^{\circ}$. However, as the azimuthal angle $\phi$ is beyond the range $(3,6)$, the difference between the observed EVPAs induced by the hairy parameter $h$ is almost independent of the observer inclination $\theta_{o}$. Furthermore, the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin $a$ than to the inclination angle $\theta_{o}$. Although the effects from the hairy parameter $h$ in the near horizon images is weaker than those from black hole spin and the observed inclination, one can infer the information on black hole’s hair by integrating the corresponding observations with higher precision, which could help to further understand the no-hair theorem.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2020YFC2201400) and National Natural Science Foundation of China (Grant Nos. 12275079 and 12035005), and the innovative research group of Hunan Province under Grant No. 2024JJ1006.

### v2

@@ III Impacts of the Hairy Parameter on the Near-Horizon Polarized Image of a rotating Horndeski Black Hole
Let us now to discuss the features of the near-horizon polarized image of a rotating Horndeski black hole (1) and to probe the corresponding impacts of the hairy parameter $h$. In Fig. 3, we present the variations of the ratio $z$ with the quantity $\Delta$ in the rotating hairy Horndeski black hole spacetime for three different plasma flow conditions as in Hou:2023bep . The plasma is assumed to infall from infinity in the type I model and to plunge from the prograde innermost stable circular orbit (ISCO) in the type II model, respectively. The type III model corresponds to the fitting result from the simulation of MAD introduced in Chael:2021rjo . Our results show that the curves $z-\Delta$ are almost overlapped in the range $\Delta<10^{-2}$ for all three types of plasma flows in the black hole spacetime with different values of the hairy parameter $h$. This further confirms that the effects arising from types of plasma flows on the polarization structures near the black hole horizon are negligible. Furthermore, from their partial enlargements, we find that the curves $z-\Delta$ are separated for different hairy parameters $h$, while still are overlapped for above three types of plasma flows. This means that the effects from the hairy parameter $h$ is much stronger than those from the types of plasma flows. In addition, the ratio $z$ increases with the increase of the hairy parameter $h$.
@@
Figure 4: Variation of the EVPA in the near-horizon images with the azimuthal angle $\phi$ in the rotating hairy Horndeski black hole spacetime. The left panel is for $\theta_{o}=17^{\circ}$ and the right one is for $\theta=80^{\circ}$. In each panel, the solid and dashed lines correspond to the cases $h=h_{c}$ and $h=0$, and the blue and red lines denote the cases $a=0.94$ and $a=0.2$, respectively.
@@
In Fig. 4, we present the changes of the observed EVPA of polarization vectors with the azimuth angle $\phi$ for the photons from region near the event horizon for a rotating hairy Horndeski black hole with different hairy parameters $h$. As expected, the observed EVPA of polarization vectors significantly changes with the black hole’s spin $a$ and the observed inclination $\theta_{o}$. Furthermore, we find that the effects of hairy parameter $h$ on the observed EVPA $\chi$ depend on the black hole’s spin $a$ and the observed inclination $\theta_{o}$. For a fixed inclination $\theta_{o}$, in the slowly rotating case where $a=0.2$, the observed EVPA $\chi$ corresponding to $h=h_{c}$ is greater than that associated with $h=0$. The situation is reversed in the rapidly rotating case with $a=0.94$. For a fixed black hole spin $a$, as the azimuthal angle $\phi$ is in the range $(3,6)$, one can find that the impact of $h$ on the observed EVPA is more distinct in the low inclination case with $\theta_{o}=17^{\circ}$. As the azimuthal angle $\phi$ is beyond the range $(3,6)$, the difference between the observed EVPAs induced by the hairy parameter $h$ is almost independent of the observer inclination $\theta_{o}$. Therefore, the observer inclination’s impact on the EVPA difference caused by the hairy parameter $h$ is also shaped by the azimuthal angle of the observed polarization vector. Through making a comparison, we also find that the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin $a$ than to the inclination angle $\theta_{o}$.
@@
Figure 5: EVPA in the near-horizon images of rotating hairy Horndeski black holes with different symbols: $\bullet$ indicates $\chi=\frac{\pi}{6}$, and $\triangle$ indicates $\chi=-\frac{\pi}{6}$. The left panel is for $a=0.2$ and the right one is for $a=0.94$. In each panel, the curves represent the near-horizon images. The solid and dashed lines correspond to the cases $h=h_{c}$ and $h=0$, and the blue and red lines denote the cases $\theta_{o}=17^{\circ}$ and $\theta_{o}=80^{\circ}$, respectively.
@@
In Fig. 5, we present the positions of the polarization vectors related to EVPA $\chi=\frac{\pi}{6}$ or $\chi=-\frac{\pi}{6}$ in the near-horizon images for a rotating hairy Horndeski black hole with spin parameters $a=0.2$ and $a=0.94$, observed at inclination angles $\theta_{o}=17^{\circ}$ and $\theta_{o}=80^{\circ}$. It is shown that the position of polarization vectors with the same EVPA $\chi=\frac{\pi}{6}$ or $\chi=-\frac{\pi}{6}$ changes with the hairy parameter $h$. Apart from the size of the event horizon image, the variation in the position of polarization vectors with the same EVPA, induced by the hairy parameter $h$, is most strongly affected by the black hole spin. Although the effects from the hairy parameter $h$ in the near-horizon images are weaker than those from the black hole spin and observed inclination, one can infer the information on the black hole’s hair by integrating the corresponding observations with the higher precision.
@@ IV Discussions and conclusions
We have investigated the observed polarization in the near-horizon images of a rotating hair black hole within the Horndeski gravity. Besides its mass and spin parameters, this black hole possesses an additional hairy parameter $h$, which originates from the Horndeski theory. Our results confirm once again that the EVPA of polarization vectors in the near-horizon images is governed exclusively by the spacetime geometry and is irrespective of the plasma flow geometry as found in Hou:2024qqo . We observe that the effects of the hairy parameter $h$ on the observed EVPA depend on the black hole’s spin $a$ and the observed inclination $\theta_{o}$. For a fixed inclination $\theta_{o}$ in the slowly rotating case, the observed EVPA corresponding to $h=h_{c}$ is greater than that associated with $h=0$. The situation is reversed in the rapidly rotating case. For a fixed black hole spin, the dependence of the difference between the observed EVPAs induced by the hairy parameter $h$ on the observer inclination $\theta_{o}$ is also influenced by the azimuthal angle of the observed polarization vector. As the azimuthal angle $\phi$ is in a specific range, for example, the range $(3,6)$ considered in this work, the impact of $h$ on the observed EVPA is more distinct in the low inclination case. However, as the azimuthal angle $\phi$ is beyond the specific range, the difference between the observed EVPAs induced by the hairy parameter $h$ is almost independent of the observer inclination $\theta_{o}$. Furthermore, the dependence of the hairy parameter’s impact on the EVPA is stronger with respect to the black hole spin $a$ than to the inclination angle $\theta_{o}$. Although the effects from the hairy parameter $h$ in the near-horizon images is weaker than those from black hole spin and the observed inclination, we can infer the information on the black hole’s hair by integrating the corresponding observations with the higher precision, which could help to further understand the no-hair theorem.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Nos. 12275079 and 12035005), the National Key Research and Development Program of China (Grant No. 2020YFC2201400), and the innovative research group of Hunan Province under Grant No. 2024JJ1006.